# When LLM-based Code Generation Meets the Software Development Process

### v1

@@ V Threats to Validity
Internal validity. Due to the generative nature of LLM, the responses may change across runs. To mitigate the threat, we try to execute the LLMs multiple times. As found in RQ1, the standard deviation of the results is small, so the generated results should be consistent. Variables such as temperature and LLM model version can have an impact on the generated code. To mitigate the threat, in RQ3, we conduct the experiments using different temperature values and model versions. We find that the temperature value only has a small effect on Pass@1, and model versions have a large impact on GPT. In RQ2, we study the impact of removing every development activity. However, having multiple development activities may have a tandem effect that further improves code quality. Future studies are needed to study the effects of different combinations of development activities in code generation.
@@
External validity. We conduct our study using two state-of-the-art benchmarks. However, as we discussed, there exist some issues in the provided tests. Moreover, the programming problems are mostly algorithmic, so the findings may not generalize to other code-generation tasks. Future studies should consider applying LCG on different and larger programming tasks. We use GPT3.5 as our underlying LLM. Although one can easily replace the LLM in our experiment, the findings may be different. Future studies on how the results of LCG change when using different LLMs.
@@
Construct validity. We try to implement an agent system that follows various software development processes. However, there are many variations of the same process model, and some variations may give better results. Future studies should further explore how changing the process models affect the code generation ability.
VI Conclusion
@@
Software development teams rely on following software process models to ensure software quality. Given that the Large Language Model (LLM) agents can act as domain experts and communicate with other agents, our study explores their efficacy in adopting engineering roles and facilitating interactions across various software process models. In this paper, we introduce LCG, a framework that implements three renowned process models: , , and . We evaluated how these models affect code generation in terms of correctness and code quality on four benchmarks: HumanEval, HumanEval-ET, MBPP, and MBPP-ET. Our findings show that notably enhances Pass@1 scores by an average of 15% over RawGPT, while maintaining the lowest standard deviation (averaging 1.2%). While GPT’s Pass@1 suffers significant variance when the model version changes (e.g., Pass@1 changes from 60 to 5 in HumanEval), the three process models still maintain stable Pass@1. Moreover, we analyzed the generated code’s quality by examining code smells and the ability to handle exceptions. Our results suggest that incorporating design and code review activities significantly reduces code smells and increases the presence of handled exceptions. This indicates that LCG not only boosts code correctness but also reduces code smells and improves reliability. These insights pave the way for future research to develop innovative development models tailored for LLM integration in software development processes.

### v2

@@ VI Threats to Validity
Internal validity. Due to the generative nature of LLM, the responses may change across runs. Variables such as temperature and LLM model version can also impact the generated code. We set the temperate value to be larger than 0 because we want LLMs to be more creative and offer more advice during discussions. To mitigate the threat, we try to execute the LLMs multiple times. As found in RQ1, the standard deviation of the results is small, so the generated results should be consistent. In RQ3, we conducted the experiments using different temperature values and model versions. The temperature value only has a small effect on Pass@1, and model versions have a large impact on RawGPT. In RQ2, we study the impact of removing every development activity. However, having multiple development activities may have a tandem effect that further improves code quality. Future studies are needed to study the effects of different combinations of development activities in code generation.
@@
External validity. We conduct our study using two state-of-the-art benchmarks. However, as we discussed, there exist some issues in the provided tests. Moreover, the programming problems are mostly algorithmic, so the findings may not generalize to other code-generation tasks. Future studies should consider applying FlowGen on different and larger programming tasks. We use GPT3.5 as our underlying LLM. Although one can easily replace the LLM in our experiment, the findings may be different. Future studies on how the results of FlowGen change when using different LLMs.
@@
Construct validity. We try to implement an agent system that follows various software development processes. However, there are many variations of the same process model, and some variations may give better results. Future studies should further explore how changing the process models affect the code generation ability. One limitation is that there is no guarantee of the correctness of the generated tests. However, we also found that the generated tests still contribute to improving the overall quality of the generated code. Similar findings are reported on CodeT [5], where generating multiple tests helps improve code correctness. However, future studies should focus on further improving the generated tests by using traditional software engineering techniques to estimate the oracles or select higher-quality tests (e.g., mutation testing).
VII Conclusion
@@
In this paper, we emulate various roles in software development, such as requirement engineers, architects, developers, and testers, using Large Language Model (LLM) agents, and structuring their interactions according to established process models. We introduce FlowGen, a framework that implements three renowned process models: , , and . We evaluated how these models affect code generation in terms of correctness and code quality on four benchmarks: HumanEval, HumanEval-ET, MBPP, and MBPP-ET. Our findings show that notably enhances Pass@1 scores by an average of 15% over RawGPT, while maintaining the lowest standard deviation (averaging 1.2%). Moreover, we find that development activities such as design and code review significantly reduce code smells and increase the presence of handled exceptions. This indicates that FlowGen not only boosts code correctness but also reduces code smells and improves reliability. Compared with other state-of-the-art techniques, and CodeT achieved similar results, with both outperforming Reflexion. Integrating CodeT into further resulted in statistically significant improvements, achieving the highest Pass@1 scores. These insights pave the way for future research to develop innovative development models tailored for LLM integration in software development processes.
In this study, we introduced FlowGen, a framework designed to emulate software process models using Large Language Model (LLM) agents, each representing roles such as requirement engineers, architects, developers, and testers. We implemented three variations of FlowGen: , , and . Our evaluation across four benchmarks–HumanEval, HumanEval-ET, MBPP, and MBPP-ET–demonstrated the superior performance of , achieving up to 31.5% improvement in Pass@1 scores over RawGPT.
Our results showed that incorporating software process models into LLM-based code generation significantly enhances code correctness, reduces code smells, and improves exception handling. FlowGenScrum consistently outperformed other models, achieving the highest Pass@1 scores and the lowest standard deviation, indicating more stable and reliable code generation.
Additionally, our comparative analysis with state-of-the-art techniques revealed that FlowGenScrum and CodeT achieved similar results, both outperforming Reflexion. Notably, integrating CodeT into FlowGenScrum resulted in statistically significant improvements, achieving the highest Pass@1 scores. This highlights the robustness and potential of combining structured software development practices with LLM capabilities.
Future research should focus on further refining these models, incorporating more sophisticated interactions, and expanding the scope of evaluation to include a broader range of software development tasks and environments. By doing so, we can better understand the capabilities and limitations of LLMs in software engineering and continue to improve their integration into practical development workflows.
These findings underscore the potential of LLMs in mimicking real-world development practices to generate higher-quality code and emphasize the importance of structured development practices in LLM-based code generation.